OpenStudy (anonymous):

f(x,y)= x^2-y^2+xy-10x+3

OpenStudy (nikvist):

what's the question?

OpenStudy (anonymous):

to find critical point and also to determent max, min or saddle point..???

OpenStudy (nikvist):

\[f(x,y)= x^2-y^2+xy-10x+3\] \[f'_x(x,y)= 2x+y-10=0\] \[f'_y(x,y)= -2y+x=0\] \[\rightarrow x=4,y=2\] \[f''_x(x,y)= 2,f''_y(x,y)= -2\]\[f''_{xy}(x,y)=f''_{yx}(x,y)=1\] \[Hessian=-2*2-1=-5\] (4,2) is saddle point